What is the electric flux through the disk of radius r = 0.1m if the uniform electric field has

a magnitude E = 2x103 N/C? What is the flux when the disk is perpendicular to the E? What is the flux when the disk is parallel to the E?


1
Expert's answer
2021-04-29T10:44:37-0400
"\Phi=\pi r^2E\cos{\alpha}"

1)


"\Phi=\pi (0.1)^2(2000)\cos{0}=62.8\ Vm"

2)


"\Phi=\pi (0.1)^2(2000)\cos{90}=0\ Vm"